OpenStudy (anonymous):

Let n represent an even integer. Which expression represents the previous even integer? A) n + 2 B) n - 2 C) n + 1 D) n - 1

OpenStudy (anonymous):

@hartnn would it be n+2?

hartnn (hartnn):

"previous" so n-2

OpenStudy (anonymous):

Let n represent an even integer. Which expression represents the next consecutive integer? A) 2n B) n - 1 C) n + 1 D) n + 2

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

next integer is n+1 next even integer is n+2

OpenStudy (anonymous):

can you please explain step by step on how did you get the answer

OpenStudy (anonymous):

Write an equation for three times a number added to 2 is 5. A) 2n + 3 = 5 B) 5n + 3 = 2 C) 3n + 2 = 5 D) 5n + 3 = 2

hartnn (hartnn):

a number : n three times a number : 3n three times a number added to 2 : 3n+2

OpenStudy (anonymous):

An integer added to 1 creates a sum of -8. What is the integer? A) 9 B) 7 C) -9 D) -7 is it b?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

-9+1 = -8

OpenStudy (perl):

n-2